OpenStudy (anonymous):

find all the cube roots of -3+3i in polar form

OpenStudy (uri):

polar form: |z| = sqrt((-3)^2+(3)^2)) the you find the argument \(\theta = y/x\)

OpenStudy (anonymous):

\[\huge -3+3i=2\sqrt{3}cis\frac{3\pi}{4}\]

OpenStudy (uri):

Polar form: \( r (cos\theta + isin\theta)\)

OpenStudy (uri):

i think he/she should solve it themselves :)

OpenStudy (anonymous):

With this level of maths, don't think it matters.

OpenStudy (anonymous):

People with this calibre of maths, should know that the answer isn't everything. They should know how they got that answer as the most important thing out of a question.

OpenStudy (anonymous):

\[\huge \sqrt[3]{-3+3i}=(2\sqrt{3})^{\frac{1}{3}}cis\frac{3\pi}{4}\times \frac{1}{3}\]

OpenStudy (anonymous):

azteck i got \[\sqrt{18}cis(3\Pi/4)\] but then i get confused with the next steps. i agree with working it out it more important than the answer i can get the answer but i cant do the working properly

OpenStudy (anonymous):

You know De Moivre's theorem correct?

OpenStudy (anonymous):

You work backwards.

OpenStudy (anonymous):

nope i dont. i have tried working it out but the notes i have dont show a good explanation of it

OpenStudy (anonymous):

When you want to find the cube of something: \[\huge (rcis\theta)^{3}=r^3cis3\theta\]

OpenStudy (anonymous):

Using that example above, what would \[\huge (rcis)^{\frac{1}{3}}=?\] Do the exact same thing as what I did but with 1/3.

OpenStudy (anonymous):

\[\huge rcis\theta\]

OpenStudy (anonymous):

sorry about that, left the theta.

OpenStudy (anonymous):

The angle is multiplied by the value of the exponent and r (the modulus) is raised by the same value of the exponent.

OpenStudy (anonymous):

\[r1/3cis1/3\]

OpenStudy (anonymous):

theta on the end

OpenStudy (anonymous):

Correct. SO using what you got so far for your question: \[\sqrt{18}cis\frac{3\pi}{4}\] You know r is \[\sqrt{18}\] and theta is\[\frac{3\pi}{4}\] What would your final result be now?

OpenStudy (anonymous):

Because: You're trying to find: \[\huge (\sqrt{18}cis\frac{3\pi}{4})^{\frac{1}{3}}\]

OpenStudy (anonymous):

\[\sqrt{18} cis (3\pi/4)?\]

OpenStudy (anonymous):

Look at my previous post. You're trying to find that.

OpenStudy (anonymous):

hangon..

OpenStudy (anonymous):

You can draw it out if the equation tool is wasting your time.

OpenStudy (anonymous):

ye.. i have no idea :/

OpenStudy (anonymous):

THere's a draw tool beside the equation tool. It could save you a tonne of time.

OpenStudy (anonymous):